OpenStudy (anonymous):

San needs to rent a car for a one-week trip in Oregon. He is considering two companies. A+ auto rental charges $175 plus $o.10 per mile. Zippy Auto rental charges $220 plus $0.05 per mile. 1. Write a equation relating the rental cost for each company to the miles driven. 2. Under what circumstances is the rental cost same for both companies? what is the cost? 3.Under what circumstances is renting from Zippy cheaper than renting from A+? 4. Suppose Sam rents a car from A+ and drives it 225 miles. What is his rental cost? I need explanation answer, help me out?

OpenStudy (anonymous):

Let x be the number of miles driven, 1. For A+ 175+0.1*x For Zippy, 220+0.05*x 2.Equate the two to find what value of x makes the two rental cost the same, 175+0.1x=220+0.05x x=900 3.For any number of miles greater than 900, Zippy works out cheaper. 4.If Sam rents it from A+, He pays, 175+0.1*225 =197.5$

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Welcome.